if cotx=5/12 anc cosx<0 what is the measure of x to the nearest degeee
Answer by lwsshak3(11628)

if cotx=5/12 anc cosx<0 what is the measure of x to the nearest degeee
cos function is negative in quadrants II and III.
cot function is positive in quadrants I and III
so, in its standard position, the terminal side of x lies in quadrant III.
Its reference angle =arc cot 5/12=arc tan 12/5=67.38º
therefore, x=180-67=113º